A Space & astronomy forum. SpaceBanter.com

If this is your first visit, be sure to check out the FAQ by clicking the link above. You may have to register before you can post: click the register link above to proceed. To start viewing messages, select the forum that you want to visit from the selection below.

Go Back   Home » SpaceBanter.com forum » Space Science » Space Station
Site Map Home Authors List Search Today's Posts Mark Forums Read Web Partners

ISS On-orbit Status, 17-07-2004



 
 
Thread Tools Display Modes
  #1  
Old July 21st 04, 02:26 PM
Jacques van Oene
external usenet poster
 
Posts: n/a
Default ISS On-orbit Status, 17-07-2004

ISS On-orbit Status 17 July 2004

All ISS systems continue to function nominally except those noted previously
or below. Saturday -- light-duty day for the ISS crew.

After breakfast, the Komandir and his Flight Engineer performed the regular
weekly 3-hr. station cleaning. ["Uborka" includes removal of food waste
products, cleaning of compartments with vacuum cleaner, wet cleaning of the
Service Module (SM) dining table and other surfaces with "Fungistat"
disinfectant and cleaning fan screens to avoid temperature rises.]

Afterwards, Padalka and Fincke held their monthly teleconference with ISS
Program Management in Houston via S-band/audio.

At 9:40am EDT the weekly planning conference (WPC) with the ground took
place, during which Mike and Gennady discussed next week's "Look-Ahead Plan"
(regularly prepared jointly by MCC-H and MCC-M planners), via S-band/audio.

For his "Saturday Science" program today, Science Officer Fincke had
selected one of the EPO (Educational Payload Operations) demos, specifically
general earth observations and pollution research. Prior to the activity,
Mike set up the video equipment for taping the session for PAO purposes.
[As for all EPO demos, the video will be used to supplement NASA educational
materials, as well as be featured at NASA education websites.]


The current run of EarthKAM (EK/Earth Knowledge Acquired by Middle School
Students) came to an end today at 10:00am when Mike Fincke powered off the
payload in the Node, then disassembled the camera equipment and returned it
to stowage. The ground uplinked thanks for the highly successful mission
and a list of participating international schools with their specific focal
areas. [The EK program is operated by a staff of undergraduate students in
the MOC (Mission Operations Center) at U. of Cal. at San Diego, using
special software that they designed to make the flight information available
to the middle school students and to process their image requests. During a
mission, the MOC is staffed 8 hours/day to process requests, communicate
with participating schools, and analyze and target the returned images.
After the mission, staff members also annotate and create captions for
images to highlight their educational applications for use in classrooms,
slide shows, and for display. With students specializing in subjects
ranging from Literature and Communications to Biology and Computer
Engineering, the ISS EK Mission Operations team brings a unique array of
talents to the program, making EK truly an outstanding feature of ISS
utilization.]

Gennady performed the periodic replenishing of the Elektron's water supply
for electrolysis, filling the KOV thermal loops' EDV container with purified
(deionized) water from the BKO multifiltration/purification column unit.
[The procedure was specially designed to prevent air bubbles from getting
into the BZh liquid unit where they could cause micropump impeller
cavitation and Elektron shutdown, as numerous past times. In the procedure,
the EDV water is drawn from the BKO and the air/liquid separator unit (GZhS)
while the crewmember checks for any air bubbles in the EDV (and, if visible,
estimates their number).]

Mike held his weekly private family conference (PFC) via S-band/audio and
Ku-band/NetMeeting video.

As per the SO's invitation, investigators of the FVMV (Fluid Merging
Viscosity Measurement) and ISSI (In-Space Soldering Investigation) payloads
uplinked requests for additional digital video clips to supplement recent
curtailed downlinks. [The downlinking is to be done at Mike's convenience
and in whatever order he prefers, with suggested priority on FVMV first,
then ISSI.]

The CDR performed the routine maintenance of the SM's SOZh environment
control & life support system, which today included the regular weekly
inspection of the BRPK air/condensate water separator system.

The crew completed their daily 2.5-hr. physical workout program on TVIS
treadmill, RED resistive exerciser and VELO ergometer with load trainer.

Last night at 9:20pm EDT, station attitude control was taken over by the
Russian MCS (motion control system) for the subsequent RCS thruster maneuver
from LVLH to XPOP attitude. Thirty minutes later control authority returned
to the US CMGs (control moment gyros). XPOP will be maintained until 7/26
(9:00am).

Weekly Science Update (Expedition Nine -- 12th):

GASMAP: Next activity will be a Routine Health Check sometime next month.

Human Research Facility/Workstation (HRF WS): Continuing.

Advanced Ultrasound (ADUM): More next month.

Hand Posture Analyzer (HPA: Nothing new.

In-Space Soldering Investigation (ISSE): The Science Officer was thanked
for his time and effort in completing the first set of experiments
supporting the ISSI last Saturday (7/10). The teleconference prior to the
experiments and communications with the SO during ISSI operations kept the
ground "well informed on what you were doing and what was occurring".
Communications were excellent and helped with the slight procedural changes
made in real time. Review of the experiment videotapes revealed melting
kinetics, wetting characteristics, and equilibrium shape attainment of the
solder charge. The accumulated liquid flux was observed to rapidly and
circumferentially translate over the surface of the molten solder ball
demonstrating, perhaps for the first time in a microgravity environment, the
"Leidenfrost" effect. Evaluation of the experimental results is expected to
promote our knowledge of fabrication and repair techniques that might be
employed during extended space exploration missions.

Pore Formation and Mobility Investigation (PFMI): Nothing new. Space
Acceleration Measurement System (SAMS): SAMS has submitted two OCRs to
accomplish the battery status modification in it's flight code.
Implementation is pending the outcome of a 7/20 ground test. The SAMS team
looks forward to returning to nominal operation in the next few weeks. SAMS
data from Increments 6,7 and 8 is being analyzed for the next PIMS increment
report. Microgravity Acceleration Measurement System (MAMS): MAMS
continues to measure the quasi-steady and vibratory microgravity
environments. A report for data collected during Increments 6,7 and 8 is
currently being written. ("Sure to be a best seller. Look for it in your
favorite bookstore soon".)

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):
Nominal.

Protein Crystal Growth Monitoring by Digital Holographic Microscope
(PromISS): Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal
Emulsions (InSPACE): Planned.

Binary Colloidal Alloy Test-3 (BCAT-3): Nothing new.

Renal Stone (RS): Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites
(SHERES): Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT): Nothing new.

Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S.
Airlock. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation
(CBOSS-FDI): Nothing new.

Biotechnology Specimen Temperature Controller (BSTC): Planned.

Yeast Group Activation Packs (Yeast GAP): Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM): The last
session for Increment 9 came to an end today (see item, above).

Earth Science Toward Exploration Research (ESTER): Nothing new.

Serial Network Flow Monitor (SNFM): The SNFM team is looking forward to the
next ADUM data opportunity.

Fluid Merging Viscosity Measurement (FMVM): The AVI files the ground has
been receiving have excellent resolution and significantly increase science
output. Thanks to Mike for working so hard to complete the operations with
four types of fluids. As it turned out, glycerin was not viscous enough to
be deployable onto the strings and was not continued, whereas honey syringes
contained crystals and were stowed. Mike's technique for drop deployments
"was excellent". Scientifically meaningful experiments were obtained with
two different viscosity silicone oils and two different viscosity corn
syrups spanning the viscosity range of 2K to 100K centi-Poise (where water
is 1 centi-Poise). The experiment operations were highly successful and VTR
video was downloaded. A procedure is being developed to heat the honey
syringes D-1 and D-2 to dissolve the crystals observed by Fincke.

Viscous Liquid Foam--Bulk Metallic Glass (Foam): Investigators are very
pleased with the results of the recent run and thanked Mike for precise
timing. The video on the third run "was fantastic. We can extract the
ampoule temperature during the run by watching the oxidation. Using that
temperature, we can set the viscosity in our foam model".

BIOPSY (Effects of Prolonged Space Flight on Human Skeletal Muscle):
Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.

Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.

Miscible Fluids in Microgravity (MFMG): Nothing new.

Educational Payload Operations (EPO): Mike Fincke was thanked for selecting
EPO for today's "Saturday Science" activity, which featured Pollution
Research. This video will help students and educators (grades 5 -8) better
understand Earth observations from the ISS.

Crew Earth Observations (CEO): Investigators are pleased to find better
light returning to ISS orbit tracks after last week's days of "dimness down
under". Unfortunately, this weekend's return to XPOP attitude will limit
good practice opportunities with the long lenses. But there is still "lots
of time" with good summer illumination ahead. Fire season is in full swing
for western North America from New Mexico to Alaska and an anticipated
active tropical weather season should begin to really ramp up by early
August. The quality of ISS/CEO imagery remains high and investigators are
particularly pleased with overlapping and logical continuity of imagery in
recent sessions at the window.

Today's CEO (crew earth observation) photo targets, limited in the current
XPOP attitude by flight rule constraints on the use of the science window,
which is available for only ~1/4 of each orbit when not facing forward (in
"ram"), were Mekong River Delta (this is a land use site where plot size and
use is rapidly changing in the new agricultural economy around Ho Chi Minh
City. Suggested was a mapping swath as near nadir along track as the crew
"reasonably could"), Internal waves, Vietnam (looking left into the Gulf of
Tonkin for any wave packets), and Patagonian Glaciers (clearer weather
persists on the warmer Pacific side of the southern Andes, so imaging small
glacier tongues should be possible. [There appears to have been a major
snowfall in southernmost Patagonia on the opposite side of the Andes.]
Looking both sides of track).


ISS Orbit (as of this morning, 8:40am EDT [= epoch]):

Mean altitude -- 359.5 km
Apogee height -- 363.7 km
Perigee height -- 355.3 km
Period -- 91.7 min.
Inclination (to Equator) -- 51.63 deg
Eccentricity -- 0.0006251
Orbits per 24-hr. day -- 15.70
Mean altitude loss in last 24 hours -- 65 m
Revolutions since FGB/Zarya launch (Nov. '98) -- 32308

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times,
see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition,
information on International Space Station sighting opportunities can be
found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human
Spaceflight website. The current location of the International Space Station
can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's
Marshall Space Flight Center.


--
---------------------------

Jacques :-)

www.spacepatches.info



Ads
 




Thread Tools
Display Modes

Posting Rules
You may not post new threads
You may not post replies
You may not post attachments
You may not edit your posts

vB code is On
Smilies are On
[IMG] code is On
HTML code is Off
Forum Jump

Similar Threads
Thread Thread Starter Forum Replies Last Post
International Space Station Status Report #40 - 2004 Jacques van Oene Space Station 0 July 18th 04 08:08 PM
International Space Station Status Report #32 - 2004 Jacques van Oene Space Station 0 June 25th 04 04:11 PM
International Space Station Status Report #29 - 2004 Jacques van Oene Space Station 0 June 5th 04 09:44 AM
International Space Station Status Report #04-27 - 2004 Jacques van Oene Space Station 0 May 25th 04 02:33 PM
Ed Lu Letter from Space #6 Jacques van Oene Space Station 0 July 4th 03 11:10 AM


All times are GMT +1. The time now is 11:47 AM.


Powered by vBulletin® Version 3.6.4
Copyright ©2000 - 2020, Jelsoft Enterprises Ltd.
Copyright 2004-2020 SpaceBanter.com.
The comments are property of their posters.